# Iwahori Matsumoto presentation for modules of Iwahori fixed functions on symmetric spaces

### v2

@@ 8 A Module over the generic Hecke algebra
In this section we will construct an abstract module, $M_{t}$, over the abstract Iwahori Hecke algebra $\mathcal{H}_{t}$, which specializes to $S(X)^{I}$ over $\mathcal{H}(G,I)$.
@@
The Iwahori hyper-graph $\Gamma_{X}$ of $X$ is defined as follows: Its vertices correspond to the $H$ orbits on $I\backslash G$. The vertex that corresponds to $x\in I\backslash G/H$ is denoted by $v_{x}$. For every face of co-dimension 1 in $\mathcal{B}_{G}$, the building of $G$, and for every $o\in\Omega$, we have a hyper edge that contains the $H$ orbits of chambers that contain the face $f$ with the color $o$.
There is an action of $W_{aff}$ on the vertices (but not on the graph), and there is a function $l_{\sigma}$ from the set of vertices to $\mathbb{Z}$, this function corresponds to the length function $l_{\sigma}$ on the orbits.
@@
Let $t$ be a variable and let $\mathcal{H}_{t}=\mathcal{H}_{t}(W_{aff},\tilde{\Delta})$ be the generic Hecke algebra of $(W_{aff},\tilde{\Delta})$ (defined in many sources, for example [Sol21]). It is defined by generators and relations as follows:
It is generated as an algebra by $t$, by $T_{s}\in\mathcal{H}_{t}(W_{aff},\tilde{\Delta})$ for any $s\in\tilde{\Delta}$ and by $T_{o}\in\mathcal{H}_{t}(W_{aff},\tilde{\Delta})$ for $o\in\Omega$. $t$ is invertable and commutes with all $T_{s}$ and $T_{o}$ and the relations that the $T_{s},T_{o}$ satisfy are:
@@
    if $s_{1},s_{2}\in\tilde{\Delta}$ $s_{1}\neq s_{2}$ and $(s_{1}s_{2})^{m}=1$ then $T_{s_{1}}T_{s_{2}}T_{s_{1}}...=T_{s_{2}}T_{s_{1}}T_{s_{2}}...$ where each product contains $m$ elements.
@@
Recall that when we specialize $t$ to be the size of the residue field $k$, $t=q=\#k$, the algebra $\mathcal{H}_{t}(W_{aff},\tilde{\Delta})/(t-q)$ is isomorphic to the Iwahori-Hecke algebra $\mathcal{H}(G,I)$.
Notice that $\mathcal{H}_{t}(W_{aff},\tilde{\Delta})/(t-1)\simeq\mathbb{C}[W_{aff}].$
The calculation of the structure constants $\gamma_{f,g}(q)$ and the fact that they depend in a polynomial way on $q$ shows that when we vary $q$ the modules $S(X)^{I}$, over the algebras $\mathcal{H}(G,I)$, admits a similar generic version.
Indeed, we show below how to use the hyper-graph $\Gamma_{X}$ in order to construct an abstract module $M_{t}=M_{t}(\Gamma_{X})$ over $\mathcal{H}_{t}(W_{aff},\tilde{\Delta})$.
@@
Let $M_{t}(\Gamma_{X})$ be the free module over $\mathbb{C}[t]$ generated by the vertices of $\Gamma_{X}$. For $s\in\tilde{\Delta}$ a simple reflection, we define an action of $T_{s}$ on $M_{t}(\Gamma_{X})$ as follows.
Let $v$ be a vertex of $\Gamma_{X}$ let $e_{v,s\times v}$ be the unique hyper-edge that contains $v$ and $s\times v$.
@@
$T_{s}v:=-v+\gamma_{v,s\times v}(t)\sum_{v^{\prime}\in e_{vx,s\times v}}v^{\prime}$
where $\gamma_{v,s\times v}(t)$ is a polynomial in $t$ determined according to theorem 7.12 with $v$ playing the role $x$, $e_{v,s\times v}$ the role of $\mathcal{O}_{f}$ and $t$ the role of $q$.
@@
that sends the generator of $M_{t}(\Gamma_{X})$ corresponding to the vertex $v_{x}$ (attached to the double coset $x\in I\backslash G/H$) to the characteristic function of $x$. Then
  • $\alpha$ induces an isomorphism $\alpha_{q}$ of vector spaces between $M_{t}(\Gamma_{X})/(t=q)$ and $S(X)^{I}$.
@@
  • The actions of $T_{s}$ and $T_{o}$ for $s\in\tilde{\Delta}$ and $o\in\Omega$ define an action of the generic Iwahori Hecke algebra $\mathcal{H}_{t}(W_{aff},\tilde{\Delta})$ on $M_{t}(\Gamma_{X})$.
@@
Thus we only need to prove the third assertion. We have to show that for any vertex $v\in\Gamma_{X}$ the relations defining $\mathcal{H}_{t}(W_{aff},\tilde{\Delta})$ hold under the actions we defined on the vertices.
Those conditions are polynomial in $t$, so it is enough to show that they hold for infinitely many values of $t$. By construction they hold for $t=q$. Let $r$ be an odd integer and let $F_{r}$ be an unramified extension of $F$ of order $r$, denote by $\mathcal{O}_{r}$ its ring of integers and by $k_{r}$ its residue field. Denote by $X(F_{r})=\mathbf{G}(F_{r})/\mathbf{H}(F_{r})$, it is enough to show that $\Gamma_{X(F_{r})}=\Gamma_{X}$ for any odd $r$.
@@
Let $\mathcal{C},\mathcal{C}^{\prime}$ be two chambers of $\mathcal{B}_{G}$, assume we have $h\in\mathbf{H}(F_{r})$ such that $h\mathcal{C}=\mathcal{C}^{\prime}$. We will find $h^{\prime}\in H$ such that $h^{\prime}\mathcal{C}=\mathcal{C}^{\prime}$ and $\omega(h)=\omega(h^{\prime})$ and this will prove step 1.
Let $\mathcal{A}\subset\mathcal{B}_{G}$ be a $\sigma$ stable apartment that contains $\mathcal{C}$ and let $\mathcal{A}^{\prime}\subset\mathcal{B}_{G}$ be a $\sigma$ stable apartment that contains $\mathcal{C}^{\prime}$. There is $g\in G$ such that $\omega(g)=\omega(h)$, $g\mathcal{C}=\mathcal{C}^{\prime}$ and $g\mathcal{A}=\mathcal{A}^{\prime}$. $g\mathcal{A}$ and $h\mathcal{A}$ are two $\sigma$ stable apartments of $\mathcal{B}_{G}^{\prime}$ that contain $g\mathcal{C}$, by proposition 3.1 we may change $h$ such that we would have $h\mathcal{A}=g\mathcal{A}$.
@@
$r$ is odd and $\mathbf{T}$ is split over $F$ so $\mathbf{T}(\mathcal{O})/\mathbf{T}^{2}(\mathcal{O})\rightarrow\mathbf{T}(\mathcal{O}_{r})/\mathbf{T}^{2}(\mathcal{O}_{r})$ is an isomorphism of finite groups. Thus we can change $g$ by an element of $\mathbf{T}(\mathcal{O})$ and assume that $g^{-1}h\in\mathbf{T}^{2}(\mathcal{O}_{r})$, so we can write $g^{-1}h=t_{1}t^{-1}_{2}$ with $t_{1},t_{2}\in\mathbf{T}^{2}(\mathcal{O}_{r})$, $\sigma(t_{1})=t^{-1}_{1}$ and $\sigma(t_{2})=t_{2}$. Changing $h$ we may assume further that $t_{2}=1$.
@@
Concluding we have $g\in G$, $h\in\mathbf{H}(F_{r})$ and $t_{1}\in\mathbf{T}^{2}(\mathcal{O}_{r})$ with $\sigma(t_{1})=t^{-1}_{1}$ and $g^{-1}h=t_{1}$. Thus we have $\sigma(g)t_{1}^{-1}=\sigma(gt_{1})=gt_{1}$ or $t^{2}_{1}=g^{-1}\sigma(g)\in G$, thus $t^{2}_{1}$ is defined over $F$. $r$ is odd therefore $t_{1}$ is also defined over $F$. Thus $h=gt_{1}$ is defined over $F$ and we are done with step 1.
@@
Step 2: There are no additional $\mathbf{H}$ orbits over $F_{r}$ which aren’t defined over $F$.
@@
The $\Omega$ colors of the chambers doesn’t depend on the field, therefore the only way a new orbit may appear is if there is a chamber in $\mathcal{B}_{G}^{\prime}$ whose $\mathbf{H}(F_{r})$ doesn’t contain any chamber in $\mathcal{B}_{G}$.
@@
Assume there is a chamber like that. There is a gallery between any two chambers of $\mathcal{B}_{G}^{\prime}$, therefore we can find $\mathcal{C},\mathcal{C}^{\prime}\in\mathcal{B}_{G}^{\prime}$ which share a face of codimension 1 such that the $\mathbf{H}(F_{r})$ orbit of $\mathcal{C}$ contains a chamber in $\mathcal{B}_{G}$, and the $\mathbf{H}(F_{r})$ orbit of $\mathcal{C}^{\prime}$ doesn’t contain a chamber in $\mathcal{B}_{G}$. Without loss of generality we may assume that $\mathcal{C}\subset\mathcal{B}_{G}$.
@@
Let $f=\mathcal{C}\cap\mathcal{C}^{\prime}$, let $\mathcal{A}$ be a $\sigma$ stable apartment of $\mathcal{B}_{G}$ that contains $\mathcal{C}$, and let $\mathcal{C}^{\prime\prime}\subset\mathcal{A}$ be a chamber that contains $f$ and different from $\mathcal{C}$. Notice that we must have $l_{\sigma}(\mathcal{C})=l_{\sigma}(\mathcal{C}^{\prime})$. Otherwise by Lemma 5.5, all chambers in $\mathcal{B}_{G}^{\prime}$ that contain $f$ will be either in the $\mathbf{H}(F_{r})$ orbit of $\mathcal{C}$ or in the $\mathbf{H}(F_{r})$ orbit of $\mathcal{C}^{\prime\prime}$. $\mathcal{C}^{\prime}$ isn’t in the $\mathbf{H}(F_{r})$ orbit of either $\mathcal{C}$ nor $\mathcal{C}^{\prime\prime}$.
@@
Let $f^{\prime}=f\cup\sigma(f)$, and let $C_{f}$ be the set of chambers in $\mathcal{B}_{G}^{\prime}$ that contain $f$. Let $G^{r}_{f^{\prime}}={\cal G}_{f^{\prime}}(k_{r})/R_{u}({\cal G}_{f^{\prime}}(k_{r}))$. By Proposition 6.6 the $\mathbf{H}(F_{r})\cap Ker(\omega)$ orbits on $C_{f}$ correspond to the $(G^{r}_{f^{\prime}})^{\sigma}$ orbits on the flag variety of $G^{r}_{f^{\prime}}$. Following the construction of this correspondences we see that it commutes with the action of $Gal(F_{r}/F)$. Checking all the cases of symmetric spaces like in proposition 7.5 we see that there are no cases where a new orbit appears after an extension of odd rank. This finishes step 2.
@@
Thus we proved that vertices of $\Gamma_{X(F_{r})}$ are the same as the vertices of $\Gamma_{X}$. The argument for step 2 also proves that the hyper edges are the same. Thus we are done. ∎
@@ 9 Examples
In this section we will demonstrate the main results of this paper for two specific examples.
@@ 9.1 $G=SL_{2},H=T$
Let $G=SL_{2}(F)$, and let $\sigma$ be conjugation by $\begin{pmatrix}-1&0\\ 0&1\end{pmatrix}$. $H=G^{\sigma}=T$ is the torus of diagonal matrices.
Let $\mathcal{B}_{G}$ be the building of $SL_{2}(F)$ which is the free $q+1$ regular graph.
Let $W_{aff}=N_{G}(T)/T^{0}$, it is the free group generated by two order 2 elements $W_{aff}=<s_{0},s_{1}>$. Denote by $l$ the length function on the Coxeter group $W_{aff}$.
@@
Let $T_{s_{0}},T_{s_{1}}$ be the generators of $H(G,I)$ that corresponds to the simple reflections $s_{0},s_{1}\in W_{aff}$.
@@
This follows from the description given by Theorem 8.5. The next proposition summarizes the results about the structure of $S(X)^{I}$ in this case. We will assume $34|q-$ 3. In the case $14|q-$ 1 the next proposition will have to be altered slightly.
@@
  • There are three orbits for the action of $T$ on the $\sigma$ stable apartments in $\mathcal{B}_{G}$. They corresponds to three different $\sigma$ stable tori.
    The representatives for $\sigma$ stable tori under the action of $T$ are given by $g^{-1}Tg$ for $g$ in $\{id,\begin{pmatrix}1&1\\ -\frac{1}{2}&\frac{1}{2}\end{pmatrix},\begin{pmatrix}1&\pi\\ -\frac{1}{2\pi}&\frac{1}{2}\end{pmatrix}\}$. Here $\pi$ is a uniformeizer of $F$.
@@
  • the length function on the orbits is, $l_{\sigma}(z_{0})=l_{\sigma}(z_{1})=0$, $l_{\sigma}(x_{w})=l_{\sigma}(y_{w})=2\lfloor\frac{l(w)}{2}\rfloor+1$.
@@
    $T_{s_{0}}z_{0}+z_{0}=T_{s_{0}}z_{1}+z_{1}=z_{0}+z_{1}+x_{s_{0}}+x_{1}$
    $T_{s_{1}}z_{0}+z_{0}=T_{s_{1}}z_{1}+z_{1}=z_{0}+z_{1}+y_{1}+y_{s_{1}}$
    $T_{s_{0}}x_{s_{0}}+x_{s_{0}}=T_{s_{0}}x_{1}+x_{1}=\frac{q-1}{2}(x_{s_{0}}+x_{1}+z_{0}+z_{1})$
    $T_{s_{1}}y_{s_{1}}+y_{s_{1}}=T_{s_{1}}y_{1}+y_{1}=\frac{q-1}{2}(y_{s_{1}}+y_{1}+z_{0}+z_{1})$
    $T_{s}x_{w}=x_{sw}\text{ if }w\notin\{s_{0},1\}\text{ or }s=s_{1}\text{ and }l(sw)>l(w)$
    $T_{s}y_{w}=y_{sw}\text{ if }w\notin\{1,s_{1}\}\text{ or }s=s_{0}\text{ and }l(sw)>l(w)$
    $T_{s}x_{w}=qx_{sw}+(q-1)x_{w}\text{ if }w\notin\{s_{0},1\}\text{ or }s=s_{1}\text{ and }l(sw)<l(w)$
    $T_{s}y_{w}=qy_{sw}+(q-1)y_{w}\text{ if }w\notin\{1,s_{1}\}\text{ or }s=s_{0}\text{ and }l(sw)<l(w)$
  • The hyper graph $\Gamma_{X}$ is the hyper graph appearing in 1 continued downwards periodically.
    Figure 1: Part of the hyper graph $\Gamma_{X}$ of the symmetric space $SL_{2}/T$. Each parallelogram represents a hyper edge with 4 vertices, each other hyper edge contains exactly 2 vertices.
Proposition 9.1 follows from Proposition 9.2 by taking the generator $z_{0}-z_{1},z_{0}+z_{1},x_{1}-x_{s_{0}}$ and $y_{1}-y_{s_{0}}$.

### v3

@@ 8 A Module over the generic Hecke algebra
In this section, we will construct an abstract module $M_{t}$ over the abstract Iwahori Hecke algebra $\mathcal{H}_{t}$, which specializes to $S(X)^{I}$ as a module over $\mathcal{H}(G,I)$.
@@
The Iwahori hyper-graph $\Gamma_{X}$ of $X$ is defined as follows: Its vertices correspond to the $H$ orbits on $G/I$. The vertex that corresponds to $x\in H\backslash G/I$ is denoted by $v_{x}$. For every facet of codimension 1 in $\mathcal{B}_{G}$, the building of $G$, and for every $o\in\Omega$, we have a hyper-edge that contains all the $H$ orbits of chambers that contain the facet $f$ with the color $o$.
There is an action of $W_{aff}$ on the vertices (but not on the graph), and there is a function $l_{\sigma}$ from the set of vertices to $\mathbb{N}$. This function corresponds to the length function $l_{\sigma}$ on the orbits.
@@
[…]
@@
    if $s_{1},s_{2}\in\tilde{\Delta}$, $s_{1}\neq s_{2}$ and $(s_{1}s_{2})^{m}=1$ then $T_{s_{1}}T_{s_{2}}T_{s_{1}}...=T_{s_{2}}T_{s_{1}}T_{s_{2}}$... where each product contains $m$ elements.
@@
When we specialize $t$ to be the size of the residue field $k$, t=q=#kt=q=\#k. The algebra $\mathcal{H}_{t}(W_{aff},\tilde{\Delta})/(t-q)$ is isomorphic to the Iwahori-Hecke algebra $\mathcal{H}(G,I)$.
Note that $\mathcal{H}_{t}(W_{aff},\tilde{\Delta})/(t-1)\simeq\mathbb{C}[W_{aff}].$
The calculation of the structure constants $\gamma_{f,g}(q)$ and the fact that they depend in a polynomial way on $q$ shows that the modules $S(X)^{I}$, over the algebras $\mathcal{H}(G,I)$, admit a similar generic version.
[…]
@@
Let $M_{t}(\Gamma_{X})$ be the free $\mathbb{C}[t]$ module generated by the vertices of $\Gamma_{X}$. For $s\in\tilde{\Delta}$ a simple reflection, we define an action of $T_{s}$ on $M_{t}(\Gamma_{X})$ as follows.
Let $v$ be a vertex of $\Gamma_{X}$. Let $e_{v,v\times s}$ be the unique hyper-edge that contains $v$ and $v\times s$.
@@
$vT_{s}=-v+\gamma_{v,v\times s}(t)\sum_{v^{\prime}\in e_{v,v\times s}}v^{\prime}$
where $\gamma_{v,v\times s}(t)$ is a polynomial in $t$ determined according to Theorem 7.13 with $v$ playing the role of $x$, $e_{v,v\times s}$ the role of $\mathcal{O}_{f}$ and $t$ the role of $q$.
@@
For every $x\in H\backslash G/I$, the map $\alpha$ sends the generator of $M_{t}(\Gamma_{X})$ corresponding to the vertex $v_{x}$ to the characteristic function of $x$. Then
  • $\alpha$ induces an isomorphism $\alpha_{q}$ of vector spaces between $M_{t}(\Gamma_{X})/(t-q)$ and $S(X)^{I}$.
@@
[…]
@@
[…]
These relations are polynomial in $t$, so it is enough to show that they hold for infinitely many values of $t$. By construction they hold for $t=q$. Let $r$ be an odd integer and let $F_{r}$ be an unramified extension of $F$ of order $r$, denote by $\mathcal{O}_{r}$ its ring of integers and by $k_{r}$ its residue field. Denote by $X(F_{r})=\mathbf{H}(F_{r})\backslash\mathbf{G}(F_{r})$, it is enough to show that $\Gamma_{X(F_{r})}=\Gamma_{X}$ for any odd $r$.
@@
Let $\mathcal{C},\mathcal{C}^{\prime}$ be two chambers of $\mathcal{B}_{G}$, assume we have $h\in\mathbf{H}(F_{r})$ such that $h\mathcal{C}=\mathcal{C}^{\prime}$. We will find $h^{\prime}\in H$ such that $h^{\prime}\mathcal{C}=\mathcal{C}^{\prime}$ and $\omega(h)=\omega(h^{\prime})$. This will prove step 1.
Let $\mathcal{A}\subset\mathcal{B}_{G}$ be a $\sigma$ stable apartment that contains $\mathcal{C}$ and let $\mathcal{A}^{\prime}\subset\mathcal{B}_{G}$ be a $\sigma$ stable apartment that contains $\mathcal{C}^{\prime}$. There is $g\in G$ such that $\omega(g)=\omega(h)$, $g\mathcal{C}=\mathcal{C}^{\prime}$ and $g\mathcal{A}=\mathcal{A}^{\prime}$. The apartments $g\mathcal{A}$ and $h\mathcal{A}$ are two $\sigma$ stable apartments of $\mathcal{B}_{G}^{\prime}$ that contain $g\mathcal{C}$. By Proposition 3.1 we may change $h$ such that we would have $h\mathcal{A}=g\mathcal{A}$.
@@
$r$ is odd and $\mathbf{T}$ is split over $F$ so $\mathbf{T}(\mathcal{O})/\mathbf{T}^{2}(\mathcal{O})\rightarrow\mathbf{T}(\mathcal{O}_{r})/\mathbf{T}^{2}(\mathcal{O}_{r})$ is an isomorphism of finite groups. Thus, we may replace $g$ by an element of $\mathbf{T}(\mathcal{O})$ and assume that $g^{-1}h\in\mathbf{T}^{2}(\mathcal{O}_{r})$, so we can write $g^{-1}h=t_{1}t^{-1}_{2}$ with $t_{1},t_{2}\in\mathbf{T}^{2}(\mathcal{O}_{r})$, $\sigma(t_{1})=t^{-1}_{1}$ and $\sigma(t_{2})=t_{2}$. Changing $h$ we may assume further that $t_{2}=1$.
@@
In conclusion, we have $g\in G$, $h\in\mathbf{H}(F_{r})$ and $t_{1}\in\mathbf{T}^{2}(\mathcal{O}_{r})$ with $\sigma(t_{1})=t^{-1}_{1}$ and $g^{-1}h=t_{1}$. We have $\sigma(g)t_{1}^{-1}=\sigma(gt_{1})=gt_{1}$ or $t^{2}_{1}=g^{-1}\sigma(g)\in G$, thus $t^{2}_{1}$ is defined over $F$. The number $r$ is odd therefore $t_{1}$ is also defined over $F$. Thus $h=gt_{1}$ is defined over $F$ and we are done with step 1.
@@
Step 2: There are no additional $\mathbf{H}$ orbits over $F_{r}$ that are not defined over $F$.
@@
The $\Omega$ colors of the chambers do not depend on the field, therefore the only way a new orbit may appear is if there is a chamber in $\mathcal{B}_{G}^{\prime}$ whose $\mathbf{H}(F_{r})$ does not contain any chamber in $\mathcal{B}_{G}$.
@@
Suppose such a chamber exists. There is a gallery between any two chambers of $\mathcal{B}_{G}^{\prime}$, therefore we can find $\mathcal{C},\mathcal{C}^{\prime}\in\mathcal{B}_{G}^{\prime}$ which share a facet of codimension 1 such that the $\mathbf{H}(F_{r})$ orbit of $\mathcal{C}$ contains a chamber in $\mathcal{B}_{G}$, and the $\mathbf{H}(F_{r})$ orbit of $\mathcal{C}^{\prime}$ does not contain a chamber in $\mathcal{B}_{G}$. Without loss of generality, we may assume that $\mathcal{C}\subset\mathcal{B}_{G}$.
@@
Let $f=\mathcal{C}\cap\mathcal{C}^{\prime}$, let $\mathcal{A}$ be a $\sigma$ stable apartment of $\mathcal{B}_{G}$ that contains $\mathcal{C}$, and let $\mathcal{C}^{\prime\prime}\subset\mathcal{A}$ be a chamber that contains $f$ and is different from $\mathcal{C}$.
We must have $l_{\sigma}(\mathcal{C})=l_{\sigma}(\mathcal{C}^{\prime})$. Otherwise by Lemma 5.5, all chambers in $\mathcal{B}_{G}^{\prime}$ that contain $f$ will be either in the $\mathbf{H}(F_{r})$ orbit of $\mathcal{C}$ or in the $\mathbf{H}(F_{r})$ orbit of $\mathcal{C}^{\prime\prime}$. But $\mathcal{C}^{\prime}$ is not in the $\mathbf{H}(F_{r})$ orbit of neither $\mathcal{C}$ nor $\mathcal{C}^{\prime\prime}$.
Let $f^{\prime}=f\cup\sigma(f)$, and let $C_{f}$ be the set of chambers in $\mathcal{B}_{G}^{\prime}$ that contain $f$.
@@
Let $G^{r}_{f^{\prime}}={\mathcal{G}}_{f^{\prime}}(k_{r})/R_{u}({\mathcal{G}}_{f^{\prime}}(k_{r}))$. By Proposition 6.6 the $\mathbf{H}(F_{r})\cap Ker(\omega)$ orbits on $C_{f}$ correspond to the $(G^{r}_{f^{\prime}})^{\sigma}$ orbits on the flag variety of $G^{r}_{f^{\prime}}$. Following the construction of this correspondence we see that it commutes with the action of $Gal(F_{r}/F)$. Checking all the cases of symmetric spaces in Proposition 7.6 we see that there are no cases where a new orbit appears after an extension of odd rank. This completes Step 2.
@@
We proved that vertices of $\Gamma_{X(F_{r})}$ are the same as the vertices of $\Gamma_{X}$. The argument for step 2 also proves that the hyper-edges are the same. Thus we are done. ∎
@@ 9 Examples
In this section, we demonstrate the main results of this paper through two specific examples.
@@ 9.1 $G=SL_{2},H=T$
Let $G=SL_{2}(F)$, and let $\sigma$ be conjugation by $\begin{pmatrix}-1&0\\ 0&1\end{pmatrix}$. Then $H=G^{\sigma}=T$ is the torus of diagonal matrices.
Let $\mathcal{B}_{G}$ be the building of $SL_{2}(F)$, which is the free $q+1$ regular graph.
Let $W_{aff}=N_{G}(T)/T^{0}$, it is the free group generated by two elements of order 2, $W_{aff}=<s_{0},s_{1}>$. Denote by $l$ the length function on the Coxeter group $W_{aff}$.
@@
Let $T_{s_{0}},T_{s_{1}}$ be the generators of $H(G,I)$ that correspond to the simple reflections $s_{0},s_{1}\in W_{aff}$.
@@
This follows from the description given by Theorem 8.5. The next proposition summarizes the results about the structure of $S(X)^{I}$ in this case. We will assume 4|q34|q-3. In the case 4|q14|q-1 the next proposition will have to be altered slightly. The difference is that in the case of 4|q34|q-3, the action of $H$ on the set of $\sigma$ stable maximal split tori has $3$ orbits. In the case of 4|q14|q-1, the same action has $5$ orbits.
@@
  • There are three orbits for the action of $T$ on the $\sigma$ stable apartments in $\mathcal{B}_{G}$. They correspond to three different $\sigma$ stable tori.
    The representatives of the $\sigma$ stable tori under the action of $T$ are given by $g^{-1}Tg$ for $g$ in $\{id,\begin{pmatrix}1&1\\ -\frac{1}{2}&\frac{1}{2}\end{pmatrix},\begin{pmatrix}1&\pi\\ -\frac{1}{2\pi}&\frac{1}{2}\end{pmatrix}\}$. Here $\pi$ is a uniformizer of $F$.
@@
  • the length function on the orbits is given by $l_{\sigma}(z_{0})=l_{\sigma}(z_{1})=0$, $l_{\sigma}(x_{w})=l_{\sigma}(y_{w})=2\lfloor\frac{l(w)}{2}\rfloor+1$.
@@
    $z_{0}T_{s_{0}}+z_{0}=z_{1}T_{s_{0}}+z_{1}=z_{0}+z_{1}+x_{s_{0}}+x_{1}$
    $z_{0}T_{s_{1}}+z_{0}=z_{1}T_{s_{1}}+z_{1}=z_{0}+z_{1}+y_{1}+y_{s_{1}}$
    $x_{s_{0}}T_{s_{0}}+x_{s_{0}}=x_{1}T_{s_{0}}+x_{1}=\frac{q-1}{2}(x_{s_{0}}+x_{1}+z_{0}+z_{1})$
    $y_{s_{1}}T_{s_{1}}+y_{s_{1}}=y_{1}T_{s_{1}}+y_{1}=\frac{q-1}{2}(y_{s_{1}}+y_{1}+z_{0}+z_{1})$
    $x_{w}T_{s}=x_{ws}\text{ if }w\notin\{s_{0},1\}\text{ or }s=s_{1}\text{ and }l(ws)>l(w)$
    $y_{w}T_{s}=y_{ws}\text{ if }w\notin\{1,s_{1}\}\text{ or }s=s_{0}\text{ and }l(ws)>l(w)$
    $x_{w}T_{s}=qx_{ws}+(q-1)x_{w}\text{ if }w\notin\{s_{0},1\}\text{ or }s=s_{1}\text{ and }l(ws)<l(w)$
    $y_{w}T_{s}=qy_{ws}+(q-1)y_{w}\text{ if }w\notin\{1,s_{1}\}\text{ or }s=s_{0}\text{ and }l(ws)<l(w)$
  • The hyper-graph $\Gamma_{X}$ is the hyper-graph appearing in Figure 1 continued downwards periodically.
    Figure 1: Part of the hyper-graph $\Gamma_{X}$ of the symmetric space $T\backslash SL_{2}$. Each parallelogram represents a hyper-edge with 4 vertices, each other hyper-edge contains exactly 2 vertices.
Proposition 9.1 follows from Proposition 9.2 by taking the generators $z_{0}-z_{1},z_{0}+z_{1},x_{1}-x_{s_{0}}$ and $y_{1}-y_{s_{0}}$.